OpenStudy (anonymous):

9. Which of the following is equal to the number sentence listed below? 13 × (10 - 6) (13 + 6) + (6 + 10) 13 × (6 + 10) 13 × 10 - 13 × 6 13 × 6 + 13 × 10

OpenStudy (anonymous):

13 X (10-6) = 13 X 4 = 52 (13 + 6) + (6 + 10) = 19 + 16 = 35 13 × (6 + 10) = 13 X 16 = 208 13 × 10 - 13 × 6 = 130 - 78 = 52 13 × 6 + 13 × 10 = 130 + 78 = 208 so third one

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw